# Detecting Collective Excitations in Self-Gravitating Bose-Einstein Condensates via Faraday Waves

### v1

@@ III.1 Stability phase diagram
where $f(\tau)$ is a periodic function with period $\pi$. The Floquet exponent $\mu$ determines the stability of solutions. When the real part of $\mu$ equals zero, the solution becomes periodic and corresponds to the boundary between stable and unstable regions. Solutions are unstable when Re$(\mu)>0$ and stable when Re$(\mu)<0$.
@@
Fig. 3 illustrates the stability phase diagrams at different Jeans frequencies. Without self-gravity ($\Omega_{J}=0$, Fig. 3(a)), instability arises exclusively from parametric driving. As $\Omega_{J}$ increases (Figs. 3(b)-(d)), the unstable region near $\lambda\approx 0$ expands significantly due to the relationship $\lambda=4\Omega_{k}^{2}/\omega^{2}$. When $\Omega_{\rm Bog}^{2}<\Omega_{J}^{2}$, $\lambda$ becomes negative and its magnitude scales as $|\lambda|\propto\Omega_{J}^{2}$, explaining the expansion of the Jeans-dominated region. Concurrently, the parametric resonance tongues shift toward larger $\lambda$ values. This shift occurs because $\lambda\propto k^{2}$ at intermediate wavevectors (where $\Omega_{k}^{2}\propto k^{2}$), and increased gravitational attraction requires shorter wavelengths to satisfy the parametric resonance condition $\lambda=\nu^{2}$, with the shift magnitude given by $\Delta\lambda\approx(4/\omega^{2})\Delta(\Omega_{J}^{2})$.
@@
Figure 3: (Color online) Stability phase diagrams under different Jeans frequencies with $\Omega_{0}/\omega=0.5$.
Fig. 4 demonstrates the interplay of Jeans instability and parametric resonance at varying driving frequencies $\omega$. At low driving frequencies ($\omega\ll\Omega_{0}$), the instability is dominated by Jeans mechanisms, manifested as a broad unstable band near $\lambda\approx 0$. As $\omega$ decreases, the parametric resonance tongues fragment and narrow significantly due to the scaling $\Delta\lambda\sim\omega^{-2}$ and $\Delta q\sim\omega^{-4}$. Ultimately, the parametric resonance features vanish completely, leaving only the expanding Jeans-dominated region. Conversely, at high $\omega$, distinct parametric resonance tongues stabilize while the Jeans region contracts. The transition between these regimes occurs near $\omega\sim\Omega_{0}$, where competition between the two mechanisms causes the observed tongue fragmentation.
Figure 4: (Color online) Evolution of stability diagrams with decreasing driving frequency $\omega$ at $\Omega_{J}/\Omega_{0}=1$, showing progressive tongue fragmentation and Jeans region expansion.
@@ III.2 The Faraday wavevector
The Faraday wavevector $k_{F}$ is derived from the instability condition $\lambda=\nu^{2}$ :
$k_{F}\xi=\sqrt{\sqrt{\frac{1}{\Omega_{0}^{2}}\left(\Omega_{J}^{2}+\frac{\nu^{2}\omega^{2}}{4}\right)+\left(\frac{\beta^{-1}}{\hbar\Omega_{0}}-1\right)^{2}}+\left(\frac{\beta^{-1}}{\hbar\Omega_{0}}-1\right)}.$ (17)
In the absence of self-gravity and thermal effects ($\Omega_{J}=0$, $\beta^{-1}=0$), this expression reduces to the result for conventional BECs Hernández-Rajkov et al. (2021).
Fig. 5 shows $k_{F}$ as a function of driving frequency $\omega$ for the first three branches. The Faraday wavevector increases with driving frequency, approaching the behavior of non-self-gravitating BECs at large $\omega$. At low driving frequencies, $k_{F}\xi$ approaches a finite value $\sqrt{\sqrt{\Omega_{J}^{2}/\Omega_{0}^{2}+1}-1}$ in SGBECs, contrasting with the $k_{F}\rightarrow 0$ behavior in non-gravitational systems. Higher $\nu$ values yield larger $k_{F}$ at fixed $\omega$, consistent with excitation of shorter-wavelength modes.
Figure 5: (Color online) Faraday wavevector versus driving frequency for $\nu=1,2,3$ with $\Omega_{J}/\Omega_{0}=1.2$ and $\beta^{-1}=0$.
Fig. 6 presents the complete phase diagram of the primary Faraday wavevector ($\nu=1$) as a function of Jeans frequency and $\beta^{-1}$. We observe a sharp increase in $k_{F}$ near $\beta^{-1}/(\hbar\Omega_{0})=1$ at low driving frequencies ($\omega\ll\Omega_{0}$), while the transition remains smooth at high frequencies ($\omega\gg\Omega_{0}$). This arises from the competition between gravitational confinement and thermal fluctuations. At $\delta\equiv\beta^{-1}/(\hbar\Omega_{0})=1$, the thermal energy $k_{B}T$ balances the mean-field interaction energy $g_{0}n$, causing a critical shift in the excitation spectrum. In the low-$\omega$ regime, gravitational dominance amplifies the square-root singularity in Eq. (17), resulting in a steep but continuous rise in $k_{F}$ for $\delta>1$. This signifies a crossover: For $\delta<1$, the system supports stable collective excitations of the SGBEC (Bogoliubov-like modes modified by self-gravity); For $\delta>1$, thermal fluctuations destabilize the condensate, triggering dissipative collapse. High-frequency driving suppresses the singularity via the $\omega^{2}$ term in Eq. (17), yielding a smooth crossover.
Figure 6: (Color online) Phase diagram of $k_{F}\xi$ versus $\beta^{-1}/(\hbar\Omega_{0})$ and $\Omega_{J}/\Omega_{0}$ for $\nu=1$.
@@ IV Numerical results
Our numerical simulations are based on solving the dimensionless damped Mathieu equation governing the perturbation amplitude $U(\tau)$ with fixed parameters $\lambda=1$ and $q=-0.1$, where $\bar{\gamma}\equiv 4\gamma/\omega$ is the dimensionless damping coefficient. This configuration satisfies the primary parametric resonance condition $2\Omega_{k}=\omega$ for $\nu=1$. Fig. 7 presents the dynamical evolution across four dissipation strengths, with each column corresponding to: (a) $\bar{\gamma}=0$ exhibiting exponential growth; (b) $\bar{\gamma}=0.01$ showing persistent instability; (c) $\bar{\gamma}=0.02$ demonstrating suppressed growth; and (d) $\bar{\gamma}=0.05$ displaying overdamped decay. The top row tracks $U(\tau)$, while the middle row shows background-subtracted density $|\Psi_{P}|^{2}-1$ to isolate Faraday waves from the uniform ground state, and the bottom row displays momentum-space density.
@@
Figure 7: (Color online) Density wave dynamics under dissipation strengths $\bar{\gamma}=0$ (a), $0.01$ (b), $0.02$ (c), $0.05$ (d) with $\lambda=1$, $q=-0.1$.
@@
For $\bar{\gamma}=0$ (Fig. 7(a)), parametric instability drives exponential growth of $U(\tau)$ with Floquet exponent $\mu\approx 0.2$, generating pronounced density waves of wavelength $\lambda_{F}=2\pi/k$ and sharp momentum-space peaks at ±$\pm k$. At $\bar{\gamma}=0.01$ (Fig. 7(b)), dissipation fails to suppress the instability. $U(\tau)$ maintains strong exponential growth with minimal amplitude reduction. This confirms that weak dissipation cannot overcome parametric resonance. At $\bar{\gamma}=0.02$ (Fig. 7(c)), the competition between dissipation and parametric driving becomes evident: while unstable evolution persists, amplitude growth is significantly suppressed, resulting in weak density modulations. Under strong dissipation ($\bar{\gamma}=0.05$, Fig. 7(d)), dissipation dominates, causing rapid decay of $U(\tau)$ within $\tau<20$ and complete suppression of density waves. This sequence illustrates the transition from parametric-resonance-dominated ($\bar{\gamma}<0.02$) to dissipation-dominated ($\bar{\gamma}>0.02$) regimes. Gravitational effects influence this competition through the $\lambda$-parameter in Eq. (12), which encodes the $\Omega_{J}$-modified excitation spectrum.
@@
In Fig. 8, we present the evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies $\Omega_{J}/\Omega_{0}=0,0.1,0.5,1.2$, with fixed parameters $\omega/\Omega_{0}=1$, $q=-0.1$, and $\bar{\gamma}=0.01$. The corresponding $\lambda$ values, determined by the relation $\lambda=1-4(\Omega_{J}/\Omega_{0})^{2}$, are $\lambda=1.00$, $0.96$, $0.00$, and $-4.76$ respectively. For no self-gravity ($\Omega_{J}/\Omega_{0}=0$, $\lambda=1$), the system resides at the center of the primary parametric resonance tongue, resulting in exponential growth of $U(\tau)$ despite weak damping ($\bar{\gamma}=0.01$), characteristic of parametric instability that seeds Faraday waves. As $\Omega_{J}/\Omega_{0}$ increases to $0.1$ ($\lambda=0.96$), the system remains within the parametric resonance tongue but approaches its boundary, exhibiting sustained growth of $U(\tau)$ at a reduced rate compared to the $\Omega_{J}=0$ case, with Faraday waves forming at diminished amplitudes due to competition between parametric driving and incipient gravitational effects. At the critical transition point ($\Omega_{J}/\Omega_{0}=0.5$, $\lambda=0$) where $\Omega_{k}=0$, parametric resonance and Jeans instability coexist, leading to irregular, non-exponential growth of $U(\tau)$ that deviates from characteristic Faraday patterns. Finally, in the Jeans-dominated regime ($\Omega_{J}/\Omega_{0}=1.2$, $\lambda=-4.76$), self-gravity overwhelms parametric driving ($\Omega_{k}^{2}<0$), causing rapid exponential growth of $U(\tau)$ characteristic of gravitational collapse that suppresses Faraday wave formation. This progression confirms the phase diagram prediction (Fig. 4) where increasing Jeans instability fragments parametric resonance structures, with the transition at $\Omega_{J}/\Omega_{0}\sim 0.5$ ($\lambda=0$) highlighting the sensitivity of driven quantum fluids to long-range attractive interactions.
@@
Figure 8: Evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies: (a) $\Omega_{J}/\Omega_{0}=0$ ($\lambda=1$), (b) $\Omega_{J}/\Omega_{0}=0.1$ ($\lambda=0.96$), (c) $\Omega_{J}/\Omega_{0}=0.5$ ($\lambda=0$), (d) $\Omega_{J}/\Omega_{0}=1.2$ ($\lambda=-4.76$). Fixed parameters: $\omega/\Omega_{0}=1$, $q=-0.1$, $\bar{\gamma}=0.01$. Exponential growth in (a) signifies parametric instability, while rapid growth in (d) reflects Jeans collapse.
@@ V Conclusion
In this work, we have demonstrated that parametric resonance induced by periodic modulation of the $s$-wave scattering length can generate Faraday waves in self-gravitating Bose-Einstein condensates (SGBECs). Through linear stability analysis of the driven Gross-Pitaevskii-Newton equations, we derived a damped Mathieu equation governing the perturbation dynamics. Floquet analysis revealed a unified stability phase diagram that captures the competitive interplay between parametric resonance instabilities and the intrinsic Jeans instability of self-gravitating systems. The Faraday wavevector was systematically analyzed, with its dependence on gravitational strength and thermal effects characterized through comprehensive phase diagrams. Numerical simulations further illustrated the formation and evolution of density wave patterns under various dissipation conditions and Jeans frequencies. The study reveals that Faraday wave patterns exhibit high sensitivity to gravitational strength, transitioning from parametric-resonance-dominated patterns to Jeans-collapse-dominated behavior as gravitational strength increases.

### v2

@@ III.1 Stability phase diagram
[…]
@@
Figure 3 illustrates the stability phase diagrams at different Jeans frequencies. Without self-gravity ($\Omega_{J}=0$, Fig. 3(a)), instability arises exclusively from parametric driving. As $\Omega_{J}$ increases (Figs. 3(b)-(d)), the unstable region near $\lambda\approx 0$ expands significantly due to the relationship $\lambda=4\Omega_{k}^{2}/\omega^{2}$. When $\Omega_{\rm Bog}^{2}<\Omega_{J}^{2}$, $\lambda$ becomes negative and its magnitude scales as $|\lambda|\propto\Omega_{J}^{2}$, explaining the expansion of the Jeans-dominated region. Concurrently, the parametric resonance tongues shift toward larger $\lambda$ values. This shift occurs because $\lambda\propto k^{2}$ at intermediate wavevectors (where $\Omega_{k}^{2}\propto k^{2}$), and increased gravitational attraction requires shorter wavelengths to satisfy the parametric resonance condition $\lambda=\nu^{2}$, with the shift magnitude given by $\Delta\lambda\approx(4/\omega^{2})\Delta(\Omega_{J}^{2})$.
@@
Figure 3: (Color online) Stability phase diagrams under different Jeans frequencies with $\Omega_{0}/\omega=0.5$. The region of Jeans instability ($\lambda<0$) is denoted by the label ”JI” in the figure.
Figure 4 illustrates the central result of our stability analysis: the driving frequency $\omega$ serves as a key control parameter governing the competition between intrinsic Jeans instability and driven parametric resonance.
Figure 4: (Color online) Stability phase diagrams for a fixed Jeans frequency $\Omega_{J}/\Omega_{0}=1$ at progressively lower driving frequencies $\omega$, showing the expansion of the Jeans-unstable region ($\lambda<0$, labeled as ”JI”) and the concomitant fragmentation and narrowing of the parametric resonance tongues. The clear transition from a parametric-resonance-dominated regime to a Jeans-dominated regime is evident.
At low driving frequencies ($\omega\ll\Omega_{0},\Omega_{J}$), the stability diagram is dominated by gravity-induced instability. The Jeans-unstable region (labeled ”JI” in Fig. 4) expands significantly, as $|\lambda|\propto\Omega_{J}^{2}/\omega^{2}$ becomes large. Consequently, the tongues associated with parametric resonance become increasingly fragmented, narrowed, and eventually overwhelmed by the expansive domain of the intrinsic Jeans instability. In this regime, the system’s dynamics are governed by gravitational collapse, and the formation of distinct Faraday wave patterns is suppressed.
Conversely, at high driving frequencies ($\omega\gg\Omega_{0},\Omega_{J}$), the influence of the Jeans instability is markedly reduced. The Jeans-unstable region (labeled ”JI”) is compressed into a narrow band near $\lambda\approx 0$ due to the scaling $\lambda\propto 1/\omega^{2}$. Simultaneously, the parametric resonance tongues shift to larger values of $\lambda$ and, crucially, become well-separated from each other and from the Jeans band. In this regime, the instability landscape is dominated by these distinct parametric resonance structures, which define clear and accessible parameter windows for the formation of Faraday waves.
The most dramatic interplay between these mechanisms occurs for driving frequencies near the characteristic system frequency ($\omega\sim\Omega_{0}$). Here, the parametric resonance tongues undergo significant fragmentation and narrowing, governed by the scalings $\Delta\lambda\sim\omega^{-2}$ and $\Delta q\sim\omega^{-4}$. This complex transition region showcases the direct competition between the two instability mechanisms.
Thus, increasing the driving frequency does not eliminate the fundamental Jeans instability but effectively suppresses its relative dominance in the parameter space. This suppression creates a well-defined regime where the driven, parametric instability becomes the dominant and observable outcome, enabling the controlled generation of Faraday waves. This tunability-from gravity-dominated collapse (as seen in the expanding ”JI” regions at low frequencies) to pattern-forming resonance-underscores the utility of periodic driving for controlling nonlinear instabilities in self-gravitating quantum fluids and reveals a rich phase diagram structure that could be exploited for precise dynamical control.
@@ III.2 The Faraday wavevector
From the instability condition $\lambda=\nu^{2}$ for the $\nu^{\text{th}}$ resonance tongue, we obtain the Faraday wavevector $k_{F}$:
$k_{F}\xi=\sqrt{\sqrt{1+\frac{1}{\Omega_{0}^{2}}\left(\Omega_{J}^{2}+\frac{\nu^{2}\omega^{2}}{4}\right)}-1}.$ (19)
In the absence of self-gravity ($\Omega_{J}=0$), this expression reduces to the well-known result for conventional BECs Hernández-Rajkov et al. (2021):
$k_{F}^{(0)}\xi=\sqrt{\sqrt{1+\left(\frac{\nu\omega}{2\Omega_{0}}\right)^{2}}-1}.$ (20)
Figure 5 shows $k_{F}$ as a function of driving frequency $\omega$ for the first three branches. The Faraday wavevector increases with driving frequency, approaching the behavior of non-self-gravitating BECs at large $\omega$. At low driving frequencies, $k_{F}\xi$ approaches a finite value $\sqrt{\sqrt{\Omega_{J}^{2}/\Omega_{0}^{2}+1}-1}$ in SGBECs, contrasting with the $k_{F}\rightarrow 0$ behavior in non-gravitational systems. Higher $\nu$ values yield larger $k_{F}$ at fixed $\omega$, consistent with excitation of shorter-wavelength modes.
Figure 5: (Color online) Faraday wavevector versus driving frequency for $\nu=1,2,3$ with $\Omega_{J}/\Omega_{0}=1.2$.
@@ IV Numerical results
Our numerical simulations are based on solving the dimensionless damped Mathieu equation governing the perturbation amplitude $U(\tau)$ with fixed parameters $\lambda=1$ and q=0.1q=-0.1, where $\bar{\gamma}\equiv 4\gamma/\omega$ is the dimensionless damping coefficient. This configuration satisfies the primary parametric resonance condition $2\Omega_{k}=\omega$ for $\nu=1$. Fig. 6 presents the dynamical evolution across four dissipation strengths, with each column corresponding to: (a) $\bar{\gamma}=0$ exhibiting exponential growth; (b) $\bar{\gamma}=0.01$ showing persistent instability; (c) $\bar{\gamma}=0.02$ demonstrating suppressed growth; and (d) $\bar{\gamma}=0.05$ displaying overdamped decay. The top row tracks $U(\tau)$, while the middle row shows background-subtracted density $|\Psi_{P}|^{2}/n_{z}-1$ to isolate Faraday waves from the uniform ground state, and the bottom row displays momentum-space density.
@@
Figure 6: (Color online) Density wave dynamics under dissipation strengths $\bar{\gamma}=0$ (a), $0.01$ (b), $0.02$ (c), $0.05$ (d) with $\lambda=1$, q=0.1q=-0.1.
@@
For $\bar{\gamma}=0$ (Fig. 6(a)), parametric instability drives exponential growth of $U(\tau)$ with Floquet exponent $\mu\approx 0.2$, generating pronounced density waves of wavelength $\lambda_{F}=2\pi/k$ and sharp momentum-space peaks at $\pm k$. At $\bar{\gamma}=0.01$ (Fig. 6(b)), dissipation fails to suppress the instability. $U(\tau)$ maintains strong exponential growth with minimal amplitude reduction. This confirms that weak dissipation cannot overcome parametric resonance. At $\bar{\gamma}=0.02$ (Fig. 6(c)), the competition between dissipation and parametric driving becomes evident: while unstable evolution persists, amplitude growth is significantly suppressed, resulting in weak density modulations. Under strong dissipation ($\bar{\gamma}=0.05$, Fig. 6(d)), dissipation dominates, causing rapid decay of $U(\tau)$ within $\tau<20$ and complete suppression of density waves. This sequence illustrates the transition from parametric-resonance-dominated ($\bar{\gamma}<0.02$) to dissipation-dominated ($\bar{\gamma}>0.02$) regimes. Gravitational effects influence this competition through the $\lambda$-parameter in Eq. (14), which encodes the $\Omega_{J}$-modified excitation spectrum.
@@
In Figure 7, we present the evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies $\Omega_{J}/\Omega_{0}=0,0.1,0.5,1.2$, with fixed parameters $\omega/\Omega_{0}=1$, q=0.1q=-0.1, and $\bar{\gamma}=0.01$. The corresponding $\lambda$ values, determined by the relation $\lambda=1-4(\Omega_{J}/\Omega_{0})^{2}$, are $\lambda=1.00$, $0.96$, $0.00$, and 4.76-4.76 respectively. For no self-gravity ($\Omega_{J}/\Omega_{0}=0$, $\lambda=1$), the system resides at the center of the primary parametric resonance tongue, resulting in exponential growth of $U(\tau)$ despite weak damping ($\bar{\gamma}=0.01$), characteristic of parametric instability that seeds Faraday waves. As $\Omega_{J}/\Omega_{0}$ increases to $0.1$ ($\lambda=0.96$), the system remains within the parametric resonance tongue but approaches its boundary, exhibiting sustained growth of $U(\tau)$ at a reduced rate compared to the $\Omega_{J}=0$ case, with Faraday waves forming at diminished amplitudes due to competition between parametric driving and incipient gravitational effects. At the critical transition point ($\Omega_{J}/\Omega_{0}=0.5$, $\lambda=0$) where $\Omega_{k}=0$, parametric resonance and Jeans instability coexist, leading to irregular, non-exponential growth of $U(\tau)$ that deviates from characteristic Faraday patterns. Finally, in the Jeans-dominated regime ($\Omega_{J}/\Omega_{0}=1.2$, $\lambda=-4.76$), self-gravity overwhelms parametric driving ($\Omega_{k}^{2}<0$), causing rapid exponential growth of $U(\tau)$ characteristic of gravitational collapse that suppresses Faraday wave formation. This progression confirms the phase diagram prediction (Fig. 4) where increasing Jeans instability fragments parametric resonance structures, with the transition at $\Omega_{J}/\Omega_{0}\sim 0.5$ ($\lambda=0$) highlighting the sensitivity of driven quantum fluids to long-range attractive interactions.
@@
Figure 7: Evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies: (a) $\Omega_{J}/\Omega_{0}=0$ ($\lambda=1$), (b) $\Omega_{J}/\Omega_{0}=0.1$ ($\lambda=0.96$), (c) $\Omega_{J}/\Omega_{0}=0.5$ ($\lambda=0$), (d) $\Omega_{J}/\Omega_{0}=1.2$ ($\lambda=-4.76$). Fixed parameters: $\omega/\Omega_{0}=1$, q=0.1q=-0.1, $\bar{\gamma}=0.01$. Exponential growth in (a) signifies parametric instability, while rapid growth in (d) reflects Jeans collapse.
The numerical evolution of the perturbation amplitude $U(\tau)$ in Fig. 7 provides a temporal signature of the competition and transition between instability mechanisms. The progression from clear oscillatory growth (parametrically driven Faraday waves) at $\lambda=1$ to pure monotonic growth (Jeans collapse) at $\lambda=-4.76$, with a crossover region of irregular growth at $\lambda=0$, corroborates the phase diagram prediction. It visually demonstrates that driving competes with the intrinsic instability, and the final outcome is determined by their relative strength in the $\lambda$-$q$ parameter space.
@@ V Conclusion
In this work, we have demonstrated that the stability landscape of a driven SGBEC is characterized by the competitive interplay between its intrinsic Jeans instability and driven parametric resonance. The central finding is that this competition can be effectively controlled and tuned via the driving frequency. High-frequency driving suppresses the prominence of Jeans instability in parameter space and enables the domination of parametric resonance, creating a well-defined regime for probing collective excitations via Faraday waves. Our work thus establishes periodic driving not as a way to eliminate fundamental instabilities, but as a tool to manipulate their competition and access novel non-equilibrium states in quantum fluids with long-range interactions.
@@ Acknowledgments
Appendix A Reduction to quasi-1D equations for self-gravitational potential
The 3D gravitational potential is given by the solution of the Poisson equation:
$\Phi(\mathbf{r},t)=-Gm\int\frac{|\Psi(\mathbf{r}^{\prime},t)|^{2}}{|\mathbf{r}-\mathbf{r}^{\prime}|}d^{3}\mathbf{r}^{\prime}.$ (21)
Substituting the factorized wave function:
$\Phi(\mathbf{r},t)=-Gm\int dz^{\prime}|\psi_{P}(z^{\prime},t)|^{2}\int d^{2}\mathbf{r}_{\perp}^{\prime}\frac{|\psi_{B}(\mathbf{r}_{\perp}^{\prime})|^{2}}{|\mathbf{r}-\mathbf{r}^{\prime}|}.$ (22)
The quasi-1D potential is defined as the transverse average:
$\Phi_{\rm 1D}(z,t)=\int|\psi_{B}(\mathbf{r}_{\perp})|^{2}\Phi(\mathbf{r},t)d^{2}\mathbf{r}_{\perp}.$ (23)
Substituting $\Phi(\mathbf{r},t)$:
$\Phi_{\rm 1D}(z,t)=-Gm\int dz^{\prime}K(z-z^{\prime})|\psi_{P}(z^{\prime},t)|^{2}.$ (24)
The integral over $\mathbf{r}_{\perp}$ and $\mathbf{r}_{\perp}^{\prime}$ defines the kernel:
$K(z-z^{\prime})=\int d^{2}\mathbf{r}_{\perp}|\psi_{B}(\mathbf{r}_{\perp})|^{2}\int d^{2}\mathbf{r}_{\perp}^{\prime}\frac{|\psi_{B}(\mathbf{r}_{\perp}^{\prime})|^{2}}{\sqrt{|\mathbf{r}_{\perp}-\mathbf{r}_{\perp}^{\prime}|^{2}+(z-z^{\prime})^{2}}}.$ (25)
Using the Fourier representation of the potential:
$\frac{1}{|\mathbf{r}|}=\frac{1}{2\pi^{2}}\int\frac{e^{i\mathbf{k}\cdot\mathbf{r}}}{k^{2}}d^{3}k,$ (26)
we can write:
$\frac{1}{\sqrt{|\mathbf{r}_{\perp}-\mathbf{r}_{\perp}^{\prime}|^{2}+(z-z^{\prime})^{2}}}=\frac{1}{2\pi^{2}}\int\frac{e^{i\mathbf{k}_{\perp}\cdot(\mathbf{r}_{\perp}-\mathbf{r}_{\perp}^{\prime})+ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}d^{2}k_{\perp}dk_{z}.$ (27)
Substituting this into $K(z-z^{\prime})$ and using the Gaussian form of $|\psi_{B}|^{2}$, we obtain:
$K(z-z^{\prime})=\frac{1}{2\pi^{2}}\int d^{2}k_{\perp}dk_{z}\frac{e^{ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}\left|\int d^{2}\mathbf{r}_{\perp}|\psi_{B}(\mathbf{r}_{\perp})|^{2}e^{i\mathbf{k}_{\perp}\cdot\mathbf{r}_{\perp}}\right|^{2}.$ (28)
The Gaussian integral evaluates to:
$\int d^{2}\mathbf{r}_{\perp}|\psi_{B}(\mathbf{r}_{\perp})|^{2}e^{i\mathbf{k}_{\perp}\cdot\mathbf{r}_{\perp}}=e^{-k_{\perp}^{2}a_{\perp}^{2}/4},$ (29)
so:
$K(z-z^{\prime})=\frac{1}{2\pi^{2}}\int d^{2}k_{\perp}dk_{z}\frac{e^{ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}e^{-k_{\perp}^{2}a_{\perp}^{2}/2}.$ (30)
Performing the $k_{z}$ integration first via contour integration:
$\int_{-\infty}^{\infty}\frac{e^{ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}dk_{z}=\frac{\pi}{k_{\perp}}e^{-k_{\perp}|z-z^{\prime}|},$ (31)
we get:
$K(z-z^{\prime})=\frac{1}{2\pi}\int_{0}^{\infty}k_{\perp}dk_{\perp}\frac{e^{-k_{\perp}|z-z^{\prime}|}}{k_{\perp}}e^{-k_{\perp}^{2}a_{\perp}^{2}/2}=\frac{1}{2\pi}\int_{0}^{\infty}e^{-k_{\perp}|z-z^{\prime}|}e^{-k_{\perp}^{2}a_{\perp}^{2}/2}dk_{\perp}.$ (32)
This integral can be evaluated exactly and yields:
$K(z-z^{\prime})=\frac{1}{\sqrt{2\pi}a_{\perp}}e^{(z-z^{\prime})^{2}/2a_{\perp}^{2}}\Gamma\left(\frac{1}{2},\frac{(z-z^{\prime})^{2}}{2a_{\perp}^{2}}\right),$ (33)
where $\Gamma(a,z)$ is the incomplete Gamma function. This expression is identically equal to:
$K(z-z^{\prime})=2\pi K_{0}\left(\frac{|z-z^{\prime}|}{\ell}\right),$ (34)
where $K_{0}$ is the modified Bessel function of the second kind and $\ell=a_{\perp}/\sqrt{2}$. Thus:
$\Phi_{\rm 1D}(z,t)=-2Gm\int_{-\infty}^{\infty}dz^{\prime}\,|\psi_{P}(z^{\prime},t)|^{2}\,K_{0}\left(\frac{|z-z^{\prime}|}{\ell}\right).$ (35)
In the long-wavelength limit ($k\ell\ll 1$), this reduces to a local form:
$\Phi_{\rm 1D}(z,t)\approx-\frac{2Gm}{a_{\perp}}|\psi_{P}(z,t)|^{2}.$ (36)
This local approximation is valid when the characteristic wavelength of density perturbations is much larger than the transverse confinement length scale $\ell$.